# Hyper-Compression: Model Compression via Hyperfunction

### v3

@@ 4.2 Algorithmic Design
For points in $[0,1]^{N}$, the core objective of the algorithm is, given the preset error $\epsilon$, to engineering as much as we can to rapidly achieve the smallest $\theta^{*}$ and the highest possible dimension $N$. First, because a large model usually has billions of parameters, searching $\theta^{*}$ for a group of $N$ parameters has to be fast to ensure that the hypercompression can scale. Second, an imprecise $\theta^{*}$ could lead to significant performance degradation in the restored networks, rendering the retraining. Therefore, the algorithm must ensure $\theta^{*}$ to be precise for the highest possible dimension $N$. Third, when time and precision allow, the algorithm should return the smallest $\theta^{*}$ for the highest compression ratio. We design a general-purpose model compression method that enjoys the following benefits (PNAS): Preferable compression ratio, No post-hoc retraining, Affordable inference time, and Short compression time, as summarized in Table 2. As Figure 5 shows, the compression is as follows:
@@
First, we flatten all parameters into a one-dimensional vector $\boldsymbol{w}=[w_{1},w_{2},\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},w_{N}]$, where $N$ is the total number of weights, and then split it into groups $\left[\boldsymbol{w}^{(1)},\boldsymbol{w}^{(2)},\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},\boldsymbol{w}^{(G)}\right]$, where $\boldsymbol{w}^{(g)}=[w_{K(g-1)+1},w_{K(g-1)+2},\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},w_{K(g-1)+K}]$ is of $\mathbb{R}^{1\times K}$, and $G$ is the number of groups. For any $\boldsymbol{w}^{(g)}$, there exists an integer $\theta^{*}_{g}$ such that $|\boldsymbol{w}^{(g)}-\tau(\theta^{*}_{g}\cdot\boldsymbol{a})|<\epsilon,g=1,\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},G$ and $\boldsymbol{a}=[a_{1},a_{2},\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},a_{K}]$. Finally, the vector $\left[\theta^{*}_{1},\theta^{*}_{2},\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},\theta^{*}_{G}\right]$ represents a compression to $\boldsymbol{w}$. Take $K=2$ and $\boldsymbol{w}=[w_{1},w_{2},w_{3},w_{4}]=[0.1,0.2,0.4,0.5]$ as an example, we split $\boldsymbol{w}$ as $[\boldsymbol{w}^{(1)},\boldsymbol{w}^{(2)}]$, where $\boldsymbol{w}^{(1)}=[w_{1},w_{2}]=[0.1,0.2]$ and $\boldsymbol{w}^{(2)}=[w_{3},w_{4}]=[0.4,0.5]$, and derive $\theta^{*}_{1}$ and $\theta^{*}_{2}$ for $\boldsymbol{w}^{(1)}$ and $\boldsymbol{w}^{(2)}$, respectively. As Figure 3 shows, we can find a $\theta$ such that $(\tau(\theta/(\pi+1)),\tau(\theta/(\pi+2)))$ can approximate a given point in the 2D space.
[…]
Table 2: The advantage comparison between hyper-compression and other model compression methods when compression ratio is high.
@@
  • In our algorithm, we set $K=2$ and replace the unit square $[0,1]^{2}$ with a more flexible box $[a,b]\times[c,d]$. Specifically, given $\boldsymbol{w}$, $[a,b]\times[c,d]$ is defined as a square with the side length $l$, centered at $(\bar{x_{i}},\bar{y_{i}})$, where $(\bar{x_{i}},\bar{y_{i}})$ is the centroid of two-dimensional points $\boldsymbol{w}^{(1)},\boldsymbol{w}^{(2)},\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},\boldsymbol{w}^{(G)}$. $l$ is a hyperparameter typically set to 0.01. Consequently, the function $\tau(x):=x-\lfloor x\rfloor$ in the ergodic theorem is generalized as $\tau([x,y]):=f([x,y])+([\bar{x_{i}},\bar{y_{i}}]-[\frac{l}{2},\frac{l}{2}])$, where $f(x)=x\bmod l$.
  • If we construct a square with its center at the centroid $(\bar{x_{i}},\bar{y_{i}})$ and the side length $l$, it is highly probable that many points in $\boldsymbol{w}^{(1)},\boldsymbol{w}^{(2)},\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},\boldsymbol{w}^{(G)}$ will fall outside this square. Therefore, we need to first scale these points into the square using a scaling factor $s$ during compression, and then scale the substituted points $\tau(\theta^{*}_{g}\cdot\boldsymbol{a})$ back during decompression. While the error rate remains intact, this inversion will amplify the error $\epsilon$, where $\epsilon=|\boldsymbol{w}^{(g)}-\tau(\theta^{*}_{g}\cdot\boldsymbol{a})|$. Assuming that $\boldsymbol{w}^{(F)}$ is the point farthest from the centroid $(\bar{x_{i}},\bar{y_{i}})$ in $\boldsymbol{w}^{(1)},\boldsymbol{w}^{(2)},\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},\boldsymbol{w}^{(G)}$, we define the farthest distance $l_{f}$ as $|\boldsymbol{w}^{(F)}-[\bar{x_{i}},\bar{y_{i}}]|$, and then the scaling factor $s=\frac{l}{2\cdot l_{f}}$. The specific formula derivation is as follows:
    $\boldsymbol{w}^{(in)}$ $=C\boldsymbol{w}^{(in)}+C$ (18)
    $=C\boldsymbol{w}^{(out)}\cdot s+C$
    $=C\boldsymbol{w}^{(out)}\cdot\frac{l}{2\cdot l_{f}}+C,$
    where $C$ is the centroid node, $\boldsymbol{w}^{(out)}$ is a node ouside the squre, and $\boldsymbol{w}^{(in)}$ is the scaled node of $\boldsymbol{w}^{(out)}$.
    Let $\boldsymbol{w}^{(in)*}=\tau(\theta^{*}_{i}\cdot\boldsymbol{a})$. Thus, $\epsilon=\boldsymbol{w}^{(in)*}-\boldsymbol{w}^{(in)}$. The detailed scale-back process is shown as follows:
    $\boldsymbol{w}^{(out)*}$ $=\frac{C\boldsymbol{w}^{(in)*}}{s}+C$ (19)
    $=\frac{\boldsymbol{w}^{(in)}+\epsilon}{s}+C.$
@@
    error $=|\boldsymbol{w}^{(out)}-\boldsymbol{w}^{(out)*}|=|C\boldsymbol{w}^{(out)}-C\boldsymbol{w}^{(out)*}|$ (20)
    $=|C\boldsymbol{w}^{(out)}-\frac{C\boldsymbol{w}^{(in)*}}{s}|=|\frac{\boldsymbol{w}^{(in)*}-\boldsymbol{w}^{(in)}}{s}|$
@@
    From (20), we can observe that the farthest distance $l_{f}$ affects the error of estimating $\boldsymbol{w}^{(out)}$ after scaling back.
    Therefore, as shown in Figure 6 in our algorithm, we classify the points outside the square into $M$ categories based on their distances from the centroid $C$, where $M$ is a hyperparameter, so that the points in different categories can utilize different farthest distance $l_{fm}$ to define different scale factor $s_{m}$. This is better than using the global farthest distance $l_{f}$ to control the compression error. This method is highly effective in practice, as most points are close to the centroid. Specifically, assuming that the points outside the square are divided into $M$ categories and $\boldsymbol{w}^{(m)}$ is a point under the $m$-th category, the scaling factor $s_{m}$ for this point is defined as
@@
    where $\lambda_{m}^{*}$ is the compression of the scaled $\boldsymbol{w}^{(m)}$ in the center square.
  • $\boldsymbol{a}=[a_{1},a_{2}]=\mathbf{d}\cdot I$ is a decomposition, where $\mathbf{d}$ and $I$ determine the direction and step size for the movement of curves defined, respectively. Since $\theta$ is an integer no more than $U$, at most we have $U+1$ points in two-dimensional space $(\tau(a_{1}\theta),\tau(a_{2}\theta)),\theta=0,1,\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},U$. In our experiment, we define an optimal vector $\boldsymbol{a}$ based on $U$, such that the distribution of $U+1$ points is more uniform, thereby minimizing the maximum error as much as possible. Specifically,
    $\boldsymbol{a}$ $=\mathbf{d}\cdot I$ (23)
@@
    where $\tan{\alpha}=\frac{l}{l/U}=U$. As shown in Figure 7, it demonstrates that using (23) results in a more uniform distribution of the sample nodes compared to defining $\boldsymbol{a}$ as follows:
    $\boldsymbol{a}=[1/(\pi+1),1/(\pi+2)].$ (24)
Figure 7: Different $\boldsymbol{a}$ will result in different distribution of points. Upper three subfigures use (23), while lower three figures utilize (24).
@@
3. KD Tree + Parallelization $\to$ Short compression time. With the advent of large models, compression time becomes an important facet of evaluating a model compression algorithm. Here, we propose a suite of techniques to enable the proposed hyper-compression method to fast finish the compression for both CPU and GPU inference.
  • When encoding the parameter vector $\boldsymbol{w}$, we first convert it into many points $\boldsymbol{w}^{(1)},\boldsymbol{w}^{(2)},\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},\boldsymbol{w}^{(G)}$. Then, given $\boldsymbol{w}^{(g)}$, we search $u$ from $\{\tau(1\cdot\boldsymbol{a}),\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},\tau(U\cdot\boldsymbol{a})\}$ to approximate $\boldsymbol{w}^{(g)}$. Next, instead of repeating searching for every point, we store $\tau(1\cdot\boldsymbol{a}),\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},\tau(U\cdot\boldsymbol{a})$ in the format of k-d trees Ram and Sinha (2019) to turn the searching problem into the problem of finding the nearest neighbor. Thus, we can fast determine $\tau(u\cdot\boldsymbol{a})\in\{\tau(1\cdot\boldsymbol{a}),\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},\tau(U\cdot\boldsymbol{a})\}$ that is closest to $\boldsymbol{w}^{(g)}$.
  • We extensively use matrix operations. Given a series of two-dimensional points $\boldsymbol{w}^{(1)},\boldsymbol{w}^{(2)},\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},\boldsymbol{w}^{(G)}$, we first calculate the scaling factor list $F$:
    $F$ $=[s_{1},s_{2},\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},s_{G}]^{\top},$ (25)
@@
    where $s_{g}\in F$ represents the scale factor of $\boldsymbol{w}^{(g)}$ and $m_{g}$ means the category of node $\boldsymbol{w}^{(g)}$. The points located within the center square are classified as category $m_{g}$.
    Let $\boldsymbol{m}=\left[m_{1},y_{2},\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},m_{G}\right]^{\top}$, $\boldsymbol{w}=\left[\boldsymbol{w}^{(1)},\boldsymbol{w}^{(2)},\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},\boldsymbol{w}^{(G)}\right]^{\top}$. Then, we can calcute the final compression $\theta^{*}$ as follows:
    $C\boldsymbol{w}=\boldsymbol{w}^{\top}-[C,\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},C]^{\top},$ (26)
    $O=C\boldsymbol{w}\cdot F+[C,\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},C]^{\top}$
    $\lambda^{*}=KD(O^{\top},[\tau(1\cdot a),\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},\tau(U\cdot\boldsymbol{a})])$
    $\theta^{*\top}=\lambda^{*\top}+U\cdot\boldsymbol{m}$
  • We also multiprocess the compression by simultaneously compressing different layers of a model to make full use of computational resources.

### v4

@@ 4.2 Algorithmic Design
For points in $0,1]^{N}$, the core objective of the algorithm is, given the preset error $\epsilon$, to engineering as much as we can to rapidly achieve the smallest $\theta^{*}$ and the highest possible dimension $N$. First, because a large model usually has billions of parameters, searching $\theta^{*}$ for a group of $N$ parameters has to be fast to ensure that the hypercompression can scale. Second, an imprecise $\theta^{*}$ could lead to significant performance degradation in the restored networks, rendering the retraining. Therefore, the algorithm must ensure $\theta^{*}$ to be precise for the highest possible dimension $N$. Third, when time and precision allow, the algorithm should return the smallest $\theta^{*}$ for the highest compression ratio. We design a general-purpose model compression method that enjoys the following benefits (PNAS): Preferable compression ratio, No post-hoc retraining, Affordable inference time, and Short compression time, as summarized in Table II. As Figure 5 shows, the compression is as follows:
@@
First, we flatten all parameters into a one-dimensional vector $\bm{w}=[w_{1},w_{2},\cdots,w_{N}]$, where $N$ is the total number of weights, and then split it into groups $\left[\bm{w}^{(1)},\bm{w}^{(2)},\cdots,\bm{w}^{(G)}\right]$, where $\bm{w}^{(g)}=[w_{K(g-1)+1},w_{K(g-1)+2},\cdots,w_{K(g-1)+K}]$ is of $\mathbb{R}^{1\times K}$, and $G$ is the number of groups. For any $\bm{w}^{(g)}$, there exists an integer $\theta^{*}_{g}$ such that $|\bm{w}^{(g)}-\tau(\theta^{*}_{g}\cdot\bm{a})|<\epsilon,g=1,\cdots,G$ and $\bm{a}=[a_{1},a_{2},\cdots,a_{K}]$. Finally, the vector $\left[\theta^{*}_{1},\theta^{*}_{2},\cdots,\theta^{*}_{G}\right]$ represents a compression to $\bm{w}$. Take $K=2$ and $\bm{w}=[w_{1},w_{2},w_{3},w_{4}]=[0.1,0.2,0.4,0.5]$ as an example, we split $\bm{w}$ as $\bm{w}^{(1)},\bm{w}^{(2)}]$, where $\bm{w}^{(1)}=[w_{1},w_{2}]=[0.1,0.2]$ and $\bm{w}^{(2)}=[w_{3},w_{4}]=[0.4,0.5]$, and derive $\theta^{*}_{1}$ and $\theta^{*}_{2}$ for $\bm{w}^{(1)}$ and $\bm{w}^{(2)}$, respectively. As Figure 3 shows, we can find a $\theta$ such that $(\tau(\theta/(\pi+1)),\tau(\theta/(\pi+2)))$ can approximate a given point in the 2D space.
Given a parameter vector $\bm{w}$, $\Theta=\left[\theta^{*}_{1},\theta^{*}_{2},\cdots,\theta^{*}_{G}\right]$ is a compression. However, we do not want $\max(\theta^{*}_{i})$ to be a large number, as this would require more storage space. Therefore, we define an integer $U$ as the upper bound of $\theta^{*}_{g}$ to ensure the preferable compression efficacy. Next, $\left[\theta^{*}_{1},\theta^{*}_{2},\cdots,\theta^{*}_{G}\right]$ can be stored using the data type uintmm, where $m=\lfloor\log_{2}(U)\rfloor+1$, which is the minimum possible number of bits without causing data overflow. In other words, given a target two-dimensional point $\bm{w}^{(g)}$ and $U$, what we are doing is finding an integer $\theta^{*}_{i}\in\left[0,U\right]$ such that the error between $\tau(\theta^{*}_{g}\cdot\bm{a})$ and $\bm{w}^{(g)}$ is minimized, where $\bm{a}=[a_{1},a_{2}]$ is the irrational direction. As $U$ increases, potentially, more integers can be explored to make the error $|\bm{w}^{(g)}-\tau(\theta^{*}_{g}\cdot\bm{a})|$ smaller. Each layer can select a different $U$, allowing important layers to choose a larger $U$ to ensure the approximation error remains sufficiently small.
TABLE II: The advantage comparison between Hyper-Compression and other model compression methods when compression ratio is high.
@@
  • In our algorithm, we set $K=2$ and replace the unit square $0,1]^{2}$ with a more flexible box $a,b]\times[c,d]$. Specifically, given $\bm{w}$, $a,b]\times[c,d]$ is defined as a square with the side length $l$, centered at $(\bar{x_{i}},\bar{y_{i}})$, where $(\bar{x_{i}},\bar{y_{i}})$ is the centroid of two-dimensional points $\bm{w}^{(1)},\bm{w}^{(2)},\cdots,\bm{w}^{(G)}$. $l$ is a hyperparameter typically set to 0.01. Consequently, the function $\tau(x):=x-\lfloor x\rfloor$ in the ergodic theorem is generalized as $\tau([x,y]):=f([x,y])+([\bar{x_{i}},\bar{y_{i}}]-[\frac{l}{2},\frac{l}{2}])$, where $f(x)=x\bmod l$.
  • If we construct a square with its center at the centroid $(\bar{x_{i}},\bar{y_{i}})$ and the side length $l$, it is highly probable that many points in $\bm{w}^{(1)},\bm{w}^{(2)},\cdots,\bm{w}^{(G)}$ will fall outside this square. Therefore, we need to first scale these points into the square using a scaling factor $s$ during compression, and then scale the substituted points $\tau(\theta^{*}_{g}\cdot\bm{a})$ back during decompression. While the error rate remains intact, this inversion will amplify the error $\epsilon$, where $\epsilon=|\bm{w}^{(g)}-\tau(\theta^{*}_{g}\cdot\bm{a})|$. Assuming that $\bm{w}^{(F)}$ is the point farthest from the centroid $(\bar{x_{i}},\bar{y_{i}})$ in $\bm{w}^{(1)},\bm{w}^{(2)},\cdots,\bm{w}^{(G)}$, we define the farthest distance $l_{f}$ as $|\bm{w}^{(F)}-[\bar{x_{i}},\bar{y_{i}}]|$, and then the scaling factor $s=\frac{l}{2\cdot l_{f}}$. The specific formula derivation is as follows:
    $\bm{w}^{(in)}$ $=C\bm{w}^{(in)}+C$ (18)
    $=C\bm{w}^{(out)}\cdot s+C$
    $=C\bm{w}^{(out)}\cdot\frac{l}{2\cdot l_{f}}+C,$
    where $C$ is the centroid node, $\bm{w}^{(out)}$ is a node ouside the squre, and $\bm{w}^{(in)}$ is the scaled node of $\bm{w}^{(out)}$.
    Let $\bm{w}^{(in)*}=\tau(\theta^{*}_{i}\cdot\bm{a})$. Thus, $\epsilon=\bm{w}^{(in)*}-\bm{w}^{(in)}$. The detailed scale-back process is shown as follows:
    $\bm{w}^{(out)*}$ $=\frac{C\bm{w}^{(in)*}}{s}+C$ (19)
    $=\frac{\bm{w}^{(in)}+\epsilon}{s}+C.$
@@
    error $=|\bm{w}^{(out)}-\bm{w}^{(out)*}|=|C\bm{w}^{(out)}-C\bm{w}^{(out)*}|$ (20)
    $=|C\bm{w}^{(out)}-\frac{C\bm{w}^{(in)*}}{s}|=|\frac{\bm{w}^{(in)*}-\bm{w}^{(in)}}{s}|$
@@
    From (20), we can observe that the farthest distance $l_{f}$ affects the error of estimating $\bm{w}^{(out)}$ after scaling back.
    Therefore, as shown in Figure 6 in our algorithm, we classify the points outside the square into $M$ categories based on their distances from the centroid $C$, where $M$ is a hyperparameter, so that the points in different categories can utilize different farthest distance $l_{fm}$ to define different scale factor $s_{m}$. This is better than using the global farthest distance $l_{f}$ to control the compression error. This method is highly effective in practice, as most points are close to the centroid.
    Specifically, assuming that the points outside the square are divided into $M$ categories and $\bm{w}^{(m)}$ is a point under the $m$-th category, the scaling factor $s_{m}$ for this point is defined as
@@
    where $\lambda_{m}^{*}$ is the compression of the scaled $\bm{w}^{(m)}$ in the center square.
  • $\bm{a}=[a_{1},a_{2}]=\mathbf{d}\cdot I$ is a decomposition, where $\mathbf{d}$ and $I$ determine the direction and step size for the movement of curves defined, respectively. Since $\theta$ is an integer no more than $U$, at most we have $U+1$ points in two-dimensional space $(\tau(a_{1}\theta),\tau(a_{2}\theta)),\theta=0,1,\cdots,U$. In our experiment, we define an optimal vector $\bm{a}$ based on $U$, such that the distribution of $U+1$ points is more uniform, thereby minimizing the maximum error as much as possible. Specifically,
    $\bm{a}$ $=\mathbf{d}\cdot I$ (23)
@@
    where $\tan{\alpha}=\frac{l}{l/U}=U$. As shown in Figure 7, it demonstrates that using (23) results in a more uniform distribution of the sample nodes compared to defining $\bm{a}$ as follows:
    $\bm{a}=[1/(\pi+1),1/(\pi+2)].$ (24)
Figure 7: Different $\bm{a}$ will result in different distribution of points. Upper three subfigures use (23), while lower three figures utilize (24).
@@
3. KD-Tree + Parallelization $\to$ Short compression time. With the advent of large models, compression time becomes an important facet of evaluating a model compression algorithm. Here, we propose a suite of techniques to enable the proposed Hyper-Compression method to fast finish the compression for both CPU and GPU inference.
  • When encoding the parameter vector $\bm{w}$, we first convert it into many points $\bm{w}^{(1)},\bm{w}^{(2)},\cdots,\bm{w}^{(G)}$. Then, given $\bm{w}^{(g)}$, we search $u^{*}$ from $\{\tau(1\cdot\bm{a}),\cdots,\tau(U\cdot\bm{a})\}$ to approximate $\bm{w}^{(g)}$. Next, instead of repeating searching for every point, we store $\tau(1\cdot\bm{a}),\cdots,\tau(U\cdot\bm{a})$ in the format of a KD-Tree [40] to turn the searching problem into the problem of finding the nearest neighbor. Thus, we can determine $\tau(u^{*}\cdot\bm{a})\in\{\tau(1\cdot\bm{a}),\cdots,\tau(U\cdot\bm{a})\}$ that is closest to $\bm{w}^{(g)}$.
  • We extensively use matrix operations. Given a series of two-dimensional points $\bm{w}^{(1)},\bm{w}^{(2)},\cdots,\bm{w}^{(G)}$, we first calculate the scaling factor list $F$:
    $F$ $=[s_{1},s_{2},\cdots,s_{G}]^{\top},$ (25)
@@
    where $s_{g}\in F$ represents the scale factor of $\bm{w}^{(g)}$ and $m_{g}$ means the category of node $\bm{w}^{(g)}$. The points located within the center square are classified as category $m_{g}$.
    Let $\bm{m}=\left[m_{1},m_{2},\cdots,m_{G}\right]^{\top}$, $\bm{w}=\left[\bm{w}^{(1)},\bm{w}^{(2)},\cdots,\bm{w}^{(G)}\right]^{\top}$. Then, we can calcute the final compression $\theta^{*}$ as follows:
    $C\bm{w}=\bm{w}^{\top}-[C,\cdots,C]^{\top},$ (26)
    $O=C\bm{w}\cdot F+[C,\cdots,C]^{\top}$
    $\bm{u}^{*}=\mathtt{KD}(O^{\top},[\tau(1\cdot a),\cdots,\tau(U\cdot\bm{a})])$
    $\theta^{*\top}=\bm{u}^{*\top}+U\cdot\bm{m},$
    where $\mathtt{KD}$ denotes executing the KD-Tree algorithm, $U$ is the length of the codebook $\tau(1\cdot\bm{a}),\cdots,\tau(U\cdot\bm{a})]$, and $\tau(\bm{u}^{*}\cdot\bm{a})$ is the closest points to the scaled points $O$.
  • When both $U$ and $G$ are large (e.g. $U=1600$ and $G=10^{9}$), the standard KD-Tree search shown in Eq.(26) remains computationally expensive due to the massive search space. To further accelerate the compression process, we optimize the standard KD-Tree, instead of directly using it as the following formula:
    $\begin{aligned} &u^{*}=\mathtt{KD}(\bm{w}^{in},[\tau(1\cdot\bm{a}),\cdots,\tau(U\cdot\bm{a})])\\ \end{aligned},$ (27)
    where $\bm{w}^{in}=[w^{in}_{1},w^{in}_{2}]$ is a single two-dimensional weight vector we need to compress.
    We observe that a full-codebook search within the KD-Tree is redundant. Since the nodes in the codebook $\tau(1\cdot\bm{a}),\cdots,\tau(U\cdot\bm{a})]$ are organized in a structured manner with continuous indices. Specifically, $U$ nodes are aligned sequentially from the bottom-left $[a,c]$ to the top-right $[b,d]$ within a bounding box $a,b]\times[c,d]$. Therefore, we can rapidly pinpoint the indices of the nodes in the two columns adjacent to a query point $\bm{w}^{(in)}$ to serve as a search subset:
[…]
    Then, we only need to perform the nearest neighbor search over a small subset of the full codebook:
    $u^{*}=\mathtt{KD}(\bm{w}^{in},[\tau(u_{l}\cdot\bm{a}),\tau((u_{l}+1)\cdot\bm{a}),\cdots,\tau(u_{r}\cdot\bm{a})]),$ (29)
    which is much faster in terms of every search.
[…]
  • We also multiprocess the compression by simultaneously compressing different layers of a model to make full use of computational resources. Regarding potential memory limits, our engineering implementation is explicitly designed to prevent out-of-memory issues when handling ultra-large models. As shown in Eq. (LABEL:pkd), the $\mathtt{KD}$ operation performs a nearest neighbor search independently. Leveraging this independence, we implement a batch processing mechanism for memory-constrained environments in parallel. Specifically, we can define a batch size to partition weights of the target model into multiple manageable subsets. The nearest neighbor search is then executed iteratively (batch-by-batch), and the partial results are finally concatenated together. This ensures that peak memory usage remains bounded, regardless of the total model size.